OpenStudy (anonymous):

Medal and Fan!!! What is the area of a regular pentagon with sides of lengths 13 and an apothem of length 8.95 in square units?

OpenStudy (anonymous):

Just add the side lingth and you should get the porimeter and that should help you from there.

Directrix (directrix):

The area of a regular polygon can be found with this formula: A = (1/2) *a*p where a is the apothem and p is the perimeter of the regular polygon.

Directrix (directrix):

A pentagon has 5 sides and a regular pentagon of side 13 has perimeter 13*5. Area = (1/2) * ( 8.95) * (13 * 5) = ? @roz_mar23

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

I guess both ways work

Directrix (directrix):

290.875 is what I got using the formula. The formula is based on the work you did.

OpenStudy (anonymous):

Ok well thank you again